# CryoFM: A Flow-based Foundation Model for Cryo-EM Densities

### v1

@@ Abstract
Cryo-electron microscopy (cryo-EM) is a powerful technique in structural biology and drug discovery, enabling the study of biomolecules at high resolution. Significant advancements by structural biologists using cryo-EM have led to the production of over 38,626 protein density maps at various resolutions [^1] The Electron Microscopy Data Bank (EMDB) is a public repository for cryo-EM volumes (https://www.ebi.ac.uk/emdb/). As of 18 September 2024, EMDB contains $38,626$ entries.. However, cryo-EM data processing algorithms have yet to fully benefit from our knowledge of biomolecular density maps, with only a few recent models being data-driven but limited to specific tasks. In this study, we present CryoFM, a foundation model designed as a generative model, learning the distribution of high-quality density maps and generalizing effectively to downstream tasks. Built on flow matching, CryoFM is trained to accurately capture the prior distribution of biomolecular density maps. Furthermore, we introduce a flow posterior sampling method that leverages CryoFM as a flexible prior for several downstream tasks in cryo-EM and cryo-electron tomography (cryo-ET) without the need for fine-tuning, achieving state-of-the-art performance on most tasks and demonstrating its potential as a foundational model for broader applications in these fields.
@@ 1 Introduction
Cryo-electron microscopy is an important technique in structural biology and drug discovery, allowing the determination of high-resolution 3D structures of biomolecules that are difficult to study through conventional methods, offering insights into molecular mechanisms and aiding in fields like drug discovery (Nogales & Scheres, 2015). A major component of cryo-EM data processing involves reconstructing 3D structures from noisy 2D projections of particles. It can be formulated as an inverse problem (Bendory et al., 2020; Singer & Sigworth, 2020), which aims at recovering the signal ${\mathbf{x}}\in\mathbb{R}^{n}$ (i.e., a clean protein density) from the observation ${\mathbf{y}}\in\mathbb{R}^{m}$  [^2] The signal ${\mathbf{x}}$ denotes a clean protein density whose shape is $\mathbb{R}^{D\times D\times D}$, where $D$ is the side length. However, the observation ${\mathbf{y}}$ is a degradation and can have any arbitary shape, so we use ${\mathbf{x}}\in\mathbb{R}^{n}$ for ease of notation but without loss of generality.. Under this formulation, following Bayesian statistics, the objective is to sample a density from the posterior $p({\mathbf{x}}|{\mathbf{y}})$, which can be factorized to $p({\mathbf{x}}|{\mathbf{y}})\propto p({\mathbf{y}}|{\mathbf{x}})p({\mathbf{x}})$. Thus, the prior distribution becomes essential for guiding the reconstruction process and improving the accuracy of the resulting structures.
@@
Scheres (2012b) first introduced a Gaussian distribution as the prior $p({\mathbf{x}})$ in 3D reconstruction for cryo-EM, effectively functioning as a frequency-dependent low-pass filter, which has proven useful in many cases. Building on this, recent works have explored more sophisticated regularizers rather than explicitly defining a prior distribution (Punjani et al., 2020; Tegunov et al., 2021; Kimanius et al., 2021; Li et al., 2023; Schwab et al., 2024; Kimanius et al., 2024; Liu et al., 2024). While these approaches were initially developed for 3D reconstruction tasks, their methodologies are closely aligned with the methods for density map modification and post-processing (Jakobi et al., 2017; Ramírez-Aportela et al., 2020; Terwilliger et al., 2020a; Kaur et al., 2021; Sanchez-Garcia et al., 2021; He et al., 2023). Both share the common objective of improving the quality of cryo-EM maps, whether during refinement or in post-processing. Among all these approaches, a line of methods has emerged that utilizes data-driven models to directly learn $p({\mathbf{x}}|{\mathbf{y}})$ from data, showing strong potential (Sanchez-Garcia et al., 2021; He et al., 2023; Kimanius et al., 2024). While powerful, these models are often designed for specific tasks, limiting their general applicability and versatility.
@@
Figure 1: The overview of CryoFM. In the training stage, CryoFM learns a vector field $v_{t}({\mathbf{x}}_{t})$, whose corresponding probability flow generates the data distribution $p_{0}({\mathbf{x}}_{0})$ of high-quality protein densities. In the inference stage, given a degraded observation ${\mathbf{y}}$, a likelihood term $p_{t}({\mathbf{y}}|{\mathbf{x}}_{t})$ is incorporated to convert the unconditional vector field $v_{t}(x_{t})$ to a conditional one $v_{t}({\mathbf{x}}_{t}|{\mathbf{y}})$, so that we can sample from the posterior distribution $p_{0}({\mathbf{x}}_{0}|{\mathbf{y}})$. This enables signal restoration of the density map, resulting in improved resolution of the alpha helices in the shown case.
@@
Foundation models leverage prior knowledge and provide versatility across a wide range of applications. In particular, they have revolutionized AI for protein structure prediction and design by utilizing vast datasets to tackle diverse tasks (Jumper et al., 2021; Baek et al., 2021; Lin et al., 2023; Ingraham et al., 2023; Krishna et al., 2024; Hayes et al., 2024; Abramson et al., 2024; Wang et al., 2024). However, experimental electron density maps, a key modality in structural biology, have largely been neglected in this space. Meanwhile, the cryo-EM field has not yet taken full advantage of foundation models, which integrate prior knowledge and offer flexibility across multiple applications. In this study, we present a foundation model, CryoFM, that directly models the prior distribution $p({\mathbf{x}})$ by learning the distribution of high-quality cryo-EM density maps through flow matching. Notably, flow matching (Lipman et al., 2022; Liu et al., 2022a) is a well-established generative model which achieves state-of-the-art results on several tasks (Dao et al., 2023; Esser et al., 2024; Yim et al., 2023a). The key ingredient of flow matching is a time-dependent vector field $v_{t}({\mathbf{x}}_{t})$, which can be employed to progressively denoise a noise vector into a data sample and is closely related to the prior term $\nabla_{{\mathbf{x}}_{t}}\log p({\mathbf{x}}_{t})$. In CryoFM, when applied to downstream tasks, we draw a sample from the posterior distribution with its score (Ho et al., 2020; Song et al., 2021b): $\nabla_{\mathbf{x}}\log p({\mathbf{x}}|{\mathbf{y}})=\nabla_{\mathbf{x}}\log p({\mathbf{x}})+\nabla_{\mathbf{x}}\log p({\mathbf{y}}|{\mathbf{x}})$. In contrast to expert-designed prior, a data-driven prior is more powerful and expressive. Compared with models that directly learn the posterior, a decomposition of the prior and likelihood is more robust and versatile. As long as we can derive the likelihood term for a downstream task, we can combine it with the learned prior in a plug-and-play fashion (Zhang et al., 2021; Venkatakrishnan et al., 2013) to sample from the posterior.
@@
CryoFM consists of two models: CryoFM-S, which captures fine local details at high resolution, and CryoFM-L, which focuses on the overall global shape of the biomolecule at medium to low resolution [^3] Unless stated otherwise, CryoFM refers to CryoFM-S throughout this paper.. Fig. 1 illustrates the overview framework of CryoFM. In the training stage, we regress the vector field with high-quality density data downloaded from the EMDB database (wwPDB Consortium, 2023). When applied to downstream tasks, we derive a flow posterior sampling method to convert the vector field $v_{t}({\mathbf{x}}_{t})$ to a conditional one $v_{t}({\mathbf{x}}_{t}|{\mathbf{y}})$, which is then able to generate a posterior distribution $p({\mathbf{x}}|{\mathbf{y}})$. When applied to various downstream tasks, CryoFM achieves state-of-the-art results on most of the experiments, demonstrating the power and potential of CryoFM as a foundation model for solving more complex problems in cryo-EM and cryo-ET.
@@
Our main contributions are as follows: (1) we present the first flow-based generative model as a foundation model that learns the distribution of high-quality cryo-EM density maps; (2) we derive a flow posterior sampling algorithm, enabling CryoFM to be effectively utilized as a prior for various downstream tasks; (3) in an unsupervised manner, we achieve better performance in several downstream tasks compared to DeepEMhancer (Sanchez-Garcia et al., 2021), EMReady (He et al., 2023) and spIsoNet (Liu et al., 2024); (4) we explore different model architectures and configurations to optimize the training of generative models on large biomolecular density maps.
@@ 2 Related Work
In this section, we will briefly review the mostly related work on cryo-EM and diffusion/flow matching for inverse problems. A more detailed discussion on these topics can be found in Section A.
@@
Diffusion/Flow Matching for Inverse Problem Inverse problems aim to restore the original sample from a degraded observation, such as in image super-resolution (Haris et al., 2018), inpainting (Yeh et al., 2017), deblurring (Kupyn et al., 2019), etc. Denoising diffusion probablistic models (DDPM) (Sohl-Dickstein et al., 2015; Ho et al., 2020) and flow-based models (Lipman et al., 2022; Liu et al., 2022a) have demonstrated their effectiveness in solving inverse problems through unsupervised approaches (Kawar et al., 2022; Chung et al., 2023b; Dao et al., 2023). Since a pretrained DDPM/Flow models the data distribution (prior distribution), it can help discover the posterior distribution given a degradation model (likelihood). Specifically, Chung et al. (2022) developed diffusion posterior sampling for general inverse problems, removing the need for strong assumptions like the linearity of the degradation operator. Another line of research examines the solution to the inverse problem when the degradation operator is unknown (Chung et al., 2023a; Kapon et al., 2024).
@@ 3.1 Flow Matching
Flow matching (Lipman et al., 2022; Liu et al., 2022a) defines the generative process of a data point ${\mathbf{x}}_{0}\in\mathbb{R}^{n}$ by progressively transforming a sample ${\mathbf{x}}_{1}\in\mathbb{R}^{n}$ from the noise distribution. The process is modeled in terms of an ordinary differential equation (ODE):
$d{\mathbf{x}}_{t}=v_{\theta}(t,{\mathbf{x}}_{t})dt,$ (1)
where $v_{\theta}:[0,1]\times\mathbb{R}^{n}\to\mathbb{R}^{n}$ is a time-dependent vector field  [^4] For ease of notaion, sometimes we omit the $\theta$ and use $v_{t}({\mathbf{x}}_{t})$ for abbreviation in later sections. parameterized by a neural network $\theta$. It denotes the velocity of a sample ${\mathbf{x}}_{t}\in\mathbb{R}^{n}$ at some intermediate timestep $t$. The vector field generates a probability density path $p_{t}({\mathbf{x}}_{t})$, which reshapes the noise distribution $p_{1}({\mathbf{x}}_{1})$ to the data distribution $p_{0}({\mathbf{x}}_{0})$. Chen et al. (2018) directly solves Eq. (1) via a differentiable neural ODE solver. Lipman et al. (2022) and Liu et al. (2022a) suggest that it is more efficient to learn a manually designed conditional vector field $u|{\mathbf{x}}_{0}:[0,1]\times\mathbb{R}^{n}\to\mathbb{R}^{n}$ given a data point ${\mathbf{x}}_{0}$. The principle of designing the vector field $u|{\mathbf{x}}_{0}$ is to simplify the generated flow trajectory. For example, the conditional flow trajectory can be a straight line between ${\mathbf{x}}_{0}$ and a noise sample ${\mathbf{x}}_{1}$:
@@
where the corresponding vector field is $u(t,{\mathbf{x}}_{t}|{\mathbf{x}}_{0})={\mathbf{x}}_{0}-{\mathbf{x}}_{1}$. Finally, the (conditional) flow matching objective becomes:
$\mathcal{L}=\|v_{\theta}(t,{\mathbf{x}}_{t})-u(t,{\mathbf{x}}_{t}|{\mathbf{x}}_{0})\|^{2}_{2}=\|v_{\theta}(t,{\mathbf{x}}_{t})-({\mathbf{x}}_{0}-{\mathbf{x}}_{1})\|^{2}_{2}.$
@@ 3.2 Diffusion Posterior Sampling
Diffusion posterior sampling (DPS) has emerged as a promising approach for solving inverse problems (Song et al., 2021b; a; Chung et al., 2023b). Given a measurement ${\mathbf{y}}\in\mathbb{R}^{m}$ derived from ${\mathbf{x}}\in\mathbb{R}^{n}$ with a degradation operator $\mathcal{A}:\mathbb{R}^{n}\to\mathbb{R}^{m}$, the goal is to sample ${\mathbf{x}}$ from the posterior $p({\mathbf{x}}|{\mathbf{y}})$. Leveraging the DDPM which learns the score of the prior distribution $\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t})$, it is straightforward to plug a likelihood term, allowing us to compute the score of the posterior:
$\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t}|{\mathbf{y}})=\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t})+\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{y}}|{\mathbf{x}}_{t}).$
However, $\log p_{t}({\mathbf{y}}|{\mathbf{x}}_{t})$ is generally intractable since the tractable likelihood $\log p_{0}({\mathbf{y}}|{\mathbf{x}}_{0})$ is only defined in the data space (i.e., $t=0$). This term requires marginalizing over all possible ${\mathbf{x}}_{0}\in p_{0}({\mathbf{x}}_{0})$. Chung et al. (2022) proposed to use a Laplace approximation of the likelihood term so that $p_{t}({\mathbf{y}}|{\mathbf{x}}_{t})\approx p_{0}({\mathbf{y}}|\hat{{\mathbf{x}}}_{0})$. The conditional score can thus be approximated by:
$\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t}|{\mathbf{y}})\approx\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t})+\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{y}}|\hat{{\mathbf{x}}}_{0}({\mathbf{x}}_{t})).$
If we assume the observation distribution $p_{0}({\mathbf{y}}|{\mathbf{x}}_{0})$ is Gaussian, taking derivative to the log-probability will produce:
$\nabla_{{\mathbf{x}}_{t}}\log p({\mathbf{y}}|{\mathbf{x}}_{t})\propto\nabla_{{\mathbf{x}}_{t}}\|{\mathbf{y}}-\mathcal{A}\hat{{\mathbf{x}}}_{0}({\mathbf{x}}_{t})\|^{2}_{2}.$
Putting everything together, we can approximate the score of the posterior distribution by:
$\nabla_{{\mathbf{x}}_{t}}\log p({\mathbf{x}}_{t}|{\mathbf{y}})\approx\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t})+\lambda_{t}\cdot\nabla_{{\mathbf{x}}_{t}}\|{\mathbf{y}}-\mathcal{A}\hat{{\mathbf{x}}}_{0}({\mathbf{x}}_{t})\|^{2}_{2},$
where $\lambda_{t}$ is a hyper-parameter which controls the step of the likelihood term.

### v2

@@ Abstract
Cryo-electron microscopy (cryo-EM) is a powerful technique in structural biology and drug discovery, enabling the study of biomolecules at high resolution. Significant advancements by structural biologists using cryo-EM have led to the production of around 40k protein density maps at various resolutions on EMDB [^1] EMDB (wwPDB Consortium, 2023b) is a public repository for cryo-EM density volumes.. However, cryo-EM data processing algorithms have yet to fully benefit from our knowledge of biomolecular density maps, with only a few recent models being data-driven but limited to specific tasks. In this study, we present CryoFM, a foundation model designed as a generative model, learning the distribution of high-quality density maps and generalizing effectively to downstream tasks. Built on flow matching, CryoFM is trained to accurately capture the prior distribution of biomolecular density maps. Furthermore, we introduce a flow posterior sampling method that leverages CryoFM as a flexible prior for several downstream tasks in cryo-EM and cryo-electron tomography (cryo-ET) without the need for fine-tuning, achieving state-of-the-art performance on most tasks and demonstrating its potential as a foundational model for broader applications in these fields.
@@ 1 Introduction
Cryo-electron microscopy is an important technique in structural biology and drug discovery, allowing the determination of high-resolution 3D structures of biomolecules that are difficult to study through conventional methods, offering insights into molecular mechanisms and aiding in fields like drug discovery (Nogales & Scheres, 2015). A major component of cryo-EM data processing involves reconstructing 3D structures from noisy 2D projections of particles. It can be formulated as an inverse problem (Bendory et al., 2020; Singer & Sigworth, 2020), which aims at recovering the signal ${\mathbf{x}}\in\mathbb{R}^{n}$ (i.e., a clean protein density) from the observation ${\mathbf{y}}\in\mathbb{R}^{m}$  [^2] The signal ${\mathbf{x}}\in\mathbb{R}^{D\times D\times D}$ denotes a clean protein density, where $D$ is the side length. The observation ${\mathbf{y}}$ can have any arbitary shape, so we use ${\mathbf{x}}\in\mathbb{R}^{n}$ for ease of notation but without loss of generality.. Under this formulation, following Bayesian statistics, the objective is to sample a density from the posterior $p({\mathbf{x}}|{\mathbf{y}})$, which can be factorized to $p({\mathbf{x}}|{\mathbf{y}})\propto p({\mathbf{y}}|{\mathbf{x}})p({\mathbf{x}})$. Thus, the prior distribution becomes essential for guiding the reconstruction process and improving the accuracy of the resulting structures.
@@
Scheres (2012b) first introduced a Gaussian distribution as the prior $p({\mathbf{x}})$ in 3D reconstruction for cryo-EM, effectively functioning as a frequency-dependent low-pass filter, which has proven useful in many cases. Building on this, recent works have explored more sophisticated regularizers rather than explicitly defining a prior distribution (Punjani et al., 2020; Tegunov et al., 2021; Kimanius et al., 2021; Li et al., 2024; Schwab et al., 2024; Kimanius et al., 2024; Liu et al., 2024). While these approaches were initially developed for 3D reconstruction tasks, their methodologies are closely aligned with the methods for density map modification and post-processing (Jakobi et al., 2017; Ramírez-Aportela et al., 2020; Terwilliger et al., 2020a; Kaur et al., 2021; Sanchez-Garcia et al., 2021; He et al., 2023). Both share the common objective of improving the quality of cryo-EM maps, whether during refinement or in post-processing. Among all these approaches, a line of methods has emerged that utilizes data-driven models to directly learn $p({\mathbf{x}}|{\mathbf{y}})$ from data, showing strong potential (Sanchez-Garcia et al., 2021; He et al., 2023; Kimanius et al., 2024). While powerful, these models are often designed for specific tasks, limiting their general applicability and versatility.
@@
Figure 1: The overview of CryoFM. In the training stage, CryoFM learns a vector field ${\textnormal{v}}_{\Theta}(t,{\mathbf{x}}_{t})$, whose corresponding probability flow generates the data distribution $p_{0}({\mathbf{x}}_{0})$ of high-quality protein densities. In the inference stage, given an observation ${\mathbf{y}}$, a likelihood term $p_{t}({\mathbf{y}}|{\mathbf{x}}_{t})$ is incorporated to convert the unconditional vector field ${\textnormal{v}}_{\Theta}(t,{\mathbf{x}}_{t})$ to a conditional one ${\textnormal{v}}_{\Theta}(t,{\mathbf{x}}_{t}|{\mathbf{y}})$, so that we can sample from the posterior distribution $p_{0}({\mathbf{x}}_{0}|{\mathbf{y}})$. This enables signal restoration of the density map, resulting in improved resolution of the alpha helices in the shown case.
@@
Foundation models leverage prior knowledge and provide versatility across a wide range of applications. In particular, they have revolutionized AI for protein structure prediction and design by utilizing vast datasets to tackle diverse tasks (Jumper et al., 2021; Baek et al., 2021; Lin et al., 2023; Krishna et al., 2024; Hayes et al., 2024; Abramson et al., 2024; Wang et al., 2024b). Chroma (Ingraham et al., 2023) and RFDiffusion (Watson et al., 2023) are two foundation models for protein structure modeling, which are broadly applicable to many structure-related tasks. However, experimental electron density maps, a key modality in structural biology, have largely been neglected in this space. Meanwhile, the cryo-EM field has not yet taken full advantage of foundation models, which integrate prior knowledge and offer flexibility across multiple applications. Wang et al. (2024a) developed a condition diffusion model in the density space, which achieves superior results by refining the map for model building, but is limited to single downstream task. In this study, we present a foundation model, CryoFM, that directly models the prior distribution $p({\mathbf{x}})$ by learning the distribution of high-quality cryo-EM density maps through flow matching. Notably, flow matching (Lipman et al., 2022; Liu et al., 2022a) is a well-established generative model, which achieves state-of-the-art results in many applications (Dao et al., 2023; Esser et al., 2024; Yim et al., 2023a). The key ingredient of flow matching is a time-dependent vector field ${\textnormal{v}}_{\Theta}(t,{\mathbf{x}}_{t})$, which can be employed to progressively denoise a noise vector into a data sample and is closely related to the score of the prior distribution $\nabla_{{\mathbf{x}}_{t}}\log p({\mathbf{x}}_{t})$. In CryoFM, when applied to downstream tasks, we draw a sample from the posterior distribution with its score (Ho et al., 2020; Song et al., 2021b): $\nabla_{\mathbf{x}}\log p({\mathbf{x}}|{\mathbf{y}})=\nabla_{\mathbf{x}}\log p({\mathbf{x}})+\nabla_{\mathbf{x}}\log p({\mathbf{y}}|{\mathbf{x}})$. In contrast to expert-designed prior, a data-driven prior is more powerful and expressive. Compared with models that directly learn the posterior, a decomposition of the prior and likelihood is more robust and versatile. As long as we can derive the likelihood term for a downstream task, we can combine it with the learned prior in a plug-and-play fashion (Zhang et al., 2021; Venkatakrishnan et al., 2013) to sample from the posterior.
@@
CryoFM consists of two models: CryoFM-S, which captures fine local details at high resolution, and CryoFM-L, which focuses on the overall global shape of the biomolecule at medium to low resolution [^3] Unless stated otherwise, CryoFM refers to CryoFM-S throughout this paper.. Fig. 1 provides an overview of CryoFM. In the training stage, we regress the vector field with high-quality density data downloaded from the EMDB database (wwPDB Consortium, 2023a). When applied to downstream tasks, we propose a flow posterior sampling method to convert the vector field ${\textnormal{v}}_{\Theta}(t,{\mathbf{x}}_{t})$ to a conditional one ${\textnormal{v}}_{\Theta}(t,{\mathbf{x}}_{t}|{\mathbf{y}})$, which can be used to sample from the posterior distribution $p({\mathbf{x}}|{\mathbf{y}})$. When applied to various downstream tasks, CryoFM achieves state-of-the-art results on most of the experiments, demonstrating the power and potential of CryoFM as a foundation model for solving more complex problems in cryo-EM and cryo-ET.
@@
Our main contributions are summarized as follows: (1) we present the first flow-based generative model as a foundation model that learns the distribution of high-quality cryo-EM density maps; (2) we derive a flow posterior sampling algorithm, enabling CryoFM to be effectively utilized as a prior for various downstream tasks; (3) in an unsupervised manner, we achieve better performance in several downstream tasks compared to DeepEMhancer (Sanchez-Garcia et al., 2021), EMReady (He et al., 2023) and spIsoNet (Liu et al., 2024); (4) we explore different model architectures and configurations to optimize the training of generative models on large biomolecular density maps.
@@ 2 Related Work
In this section, we will briefly review the mostly related work on cryo-EM and diffusion/flow matching for inverse problems. A more detailed discussion on these topics can be found in Section B.
@@
Diffusion/Flow Matching for Inverse Problem Inverse problems refer to a class of problems where the goal is to recover the original sample from the forward model’s observation, such as image super-resolution (Haris et al., 2018), inpainting (Yeh et al., 2017), deblurring (Kupyn et al., 2019), etc. Since a pretrained DDPM/Flow models the data distribution (prior distribution), it can help discover the posterior distribution given a forward model (likelihood). Specifically, Chung et al. (2022) developed diffusion posterior sampling for general inverse problems, removing the need for strong assumptions like the linearity of the forward operator. Another line of research examines the solution to the inverse problem when the forward operator is unknown (Chung et al., 2023a; Kapon et al., 2024). For protein structure modeling, Levy et al. (2024) adopts Chroma (Ingraham et al., 2023) as the diffusion prior for a dozen of downstream tasks, including atomic model refinement in cryo-EM.
@@ 3 Preliminaries
In this section, we briefly review two key components of cryoFM. In the training stage, flow matching (Section 3.1) is employed as the framework to learn the distribution $p(\mathrm{density})$. In the inference stage, given some observations (noisy densities, 2D projections, etc.), we use the posterior sampling (Section 3.2) technique to sample a density from $p(\mathrm{density}|\mathrm{observation})$.
@@ 3.1 Flow Matching
Flow matching (Lipman et al., 2022; Liu et al., 2022a) defines the generative process of a data point ${\mathbf{x}}_{0}\in\mathbb{R}^{n}$ by continuously transforming a sample ${\mathbf{x}}_{1}\in\mathbb{R}^{n}$ from the noise distribution, where the state at an intermediate timestep $t\in[0,1]$ is denoted by ${\mathbf{x}}_{t}\in\mathbb{R}^{n}$. A flow is the path traversed by the state ${\mathbf{x}}_{t}$, which is characterized by an ordinary differential equation (ODE):
$d{\mathbf{x}}_{t}={\textnormal{v}}_{\Theta}(t,{\mathbf{x}}_{t})dt,$ (1)
[…]
Given a set of data points sampled from $p_{1}({\mathbf{x}}_{1})$ and $p_{0}({\mathbf{x}}_{0})$, flow matching aims to learn the vector field ${\textnormal{v}}_{\Theta}$ from the data. Chen et al. (2018) directly solves Eq. (1) via a differentiable neural ODE solver. Lipman et al. (2022) and Liu et al. (2022a) suggest that it is more efficient to learn a manually designed conditional vector field ${\textnormal{u}}_{t}(\cdot|{\mathbf{x}}_{0}):[0,1]\times\mathbb{R}^{n}\to\mathbb{R}^{n}$ given a data point ${\mathbf{x}}_{0}$. For example, the conditional flow can be a straight line between ${\mathbf{x}}_{0}$ and a noise sample ${\mathbf{x}}_{1}$:
@@
where the corresponding vector field is ${\textnormal{u}}_{t}({\mathbf{x}}_{t}|{\mathbf{x}}_{0})={\mathbf{x}}_{1}-{\mathbf{x}}_{0}$. Finally, the conditional flow matching objective (Lipman et al., 2022) is defined as follows:
$\mathcal{L}(\Theta)=\mathbb{E}_{t\sim\mathcal{U}[0,1],{\mathbf{x}}_{0}\sim p_{0}({\mathbf{x}}_{0}),{\mathbf{x}}_{1}\sim p_{1}({\mathbf{x}}_{1})}\|{\textnormal{v}}_{\Theta}(t,{\mathbf{x}}_{t})-({\mathbf{x}}_{1}-{\mathbf{x}}_{0})\|^{2}_{2},$
which is equivalent to optimizing the original flow matching objective (Lipman et al., 2022).
@@ 3.2 Diffusion Posterior Sampling
Diffusion posterior sampling (DPS) has emerged as a promising approach for solving inverse problems (Song et al., 2021b; a; Chung et al., 2023b). Given a measurement ${\mathbf{y}}\in\mathbb{R}^{m}$ derived from ${\mathbf{x}}\in\mathbb{R}^{n}$ with a forward operator $\mathcal{A}:\mathbb{R}^{n}\to\mathbb{R}^{m}$, the goal is to sample ${\mathbf{x}}$ from the posterior $p({\mathbf{x}}|{\mathbf{y}})$. By Bayes’ theorem, it is easy to show that the score of the posterior distribution is:
$\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t}|{\mathbf{y}})=\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t})+\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{y}}|{\mathbf{x}}_{t}).$ (3)
[…]
$\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t}|{\mathbf{y}})\approx\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t})+\nabla_{{\mathbf{x}}_{t}}\log p_{0}({\mathbf{y}}|\hat{{\mathbf{x}}}_{0}({\mathbf{x}}_{t})),$
If we assume the observation distribution $p_{0}({\mathbf{y}}|{\mathbf{x}})$ is Gaussian conditioned on ${\mathbf{x}}$, we can show that the gradient of the log-likelihood satisfies:
$\nabla_{{\mathbf{x}}_{t}}\log p_{0}({\mathbf{y}}|\hat{{\mathbf{x}}}_{0}({\mathbf{x}}_{t}))\propto-\nabla_{{\mathbf{x}}_{t}}\|{\mathbf{y}}-\mathcal{A}\hat{{\mathbf{x}}}_{0}({\mathbf{x}}_{t})\|^{2}_{2}.$ (4)
Putting all these things together, we can approximate the score of the posterior distribution by:
$\nabla_{{\mathbf{x}}_{t}}\log p({\mathbf{x}}_{t}|{\mathbf{y}})\approx\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t})-\lambda_{t}\cdot\nabla_{{\mathbf{x}}_{t}}\|{\mathbf{y}}-\mathcal{A}\hat{{\mathbf{x}}}_{0}({\mathbf{x}}_{t})\|^{2}_{2},$
where $\lambda_{t}$ is associated with the partition function of Gaussian, and we treat it as a hyperparameter to control the step size of the likelihood term.